# biblatexref.bib

### v2

@@ 3. Scroll method for homogeneous varieties
In this section, we work out the scroll method introduced by Coskun and Riedl in [CR19] to a more general setting. This technique builds a surface scroll on the curve $Y_{t}$ whose degree can be used to improve the bound for algebraic hyperbolicity obtained in Proposition 2.9.
Here we will work on homogeneous varieties satisfying the Main Setting 1.3 which we recall here:
Let $A$ be a smooth complex variety of dimension $D$ such that:
@@
Let $\mathcal{E}$ be the line bundle $d_{1}H_{1}+\cdots+d_{m}H_{m}$ for $d_{1},\ldots,d_{m}>0$, invariant under $G$, and assume that $H_{1},\ldots,H_{m}$ is a collection of section-dominating line bundles for $\mathcal{E}$. To simplify the notation, we write $X$ for the zero locus of a very general section of $\mathcal{E}$, and $C$ for the curve in $X$, instead of $X_{t}$ and $Y_{t}$.
@@
are $\mathbb{P}^{N_{1}}$. We define a $\mathbb{P}^{N_{1}}$-line as a line contained in one of these fibers. Equivalently, a $\mathbb{P}^{N_{1}}$-line is a curve of numerical class $H_{1}^{N_{1}-1}H_{2}^{N_{2}}\cdots H_{m}^{N_{m}}$. We define $\mathbb{P}^{N_{i}}$-lines similarly for $i=1,\ldots,m$. We call a surface scroll a $\mathbb{P}^{N_{i}}$-scroll if there is a $\mathbb{P}^{N_{i}}$-line through every point of it.
@@
Let $L$ be a $\mathbb{P}^{N_{i}}$-line in $A$. Since $A$ is a homogeneous variety, its tangent bundle $T_{A}$ is globally generated, thus $T_{A}|_{L}$ is globally generated. A classic deformation argument then shows that $\dim F_{1}^{N_{i}}(A)=h^{0}(L,N_{L/A})=\dim A-K_{A}\cdot L-3=\dim A-a_{i}-3$ (cf [HDAG, Chapter 2].) ∎
@@
(cf [CR19, Lemma 2.13]) A rank one quotient $Q$ of $M_{H_{i}}|_{C}$ induces a $\mathbb{P}^{N_{i}}$-scroll $\Sigma$ over $C$ of $H_{i}$-degree equal to $\deg Q+(H_{i}\cdot C)$.
@@
Without loss of generality, assume $s_{1}\geq\cdots\geq s_{m}$. As before, we can assume $s_{1}+\cdots+s_{m}\leq\operatorname{rk}N_{C/X}=D-2$. Assume also that $s_{1}$ is the minimum possible for which the map $\beta$ is surjective. We divide the proof into the cases $s_{1}\leq D-3$, and the case $s_{1}=D-2$.
@@
Case 2. Suppose now that $s_{1}=D-2$, so we have $s_{2}=\cdots=s_{m}=0$. In this case, we have a surjection
@@
Since we are assuming this is the minimum value for $s_{1}$ for which $\beta$ is surjective, this means that the induced map
@@
has a rank-one image $Q$ (cf [Cl03].) By Lemma 3.4, this induces a $\mathbb{P}^{N_{1}}$-scroll $\Sigma\subset\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$ containing $C$ of $H_{1}$-degree $\Sigma\cdot H_{1}^{2}=\deg Q+(H_{1}\cdot C)$. And since $\Sigma$ is a $\mathbb{P}^{N_{1}}$-scroll, we have $H_{1}H_{j}\cdot\Sigma=H_{j}\cdot C$ for $j\neq 1$. Thus, $\Sigma$ has numerical class
@@
Suppose first that $\Sigma\subset Z$. We will prove in Theorem 4.2 that $X$ does not contain $\mathbb{P}^{N_{1}}$-lines. Thus, we must have $\Sigma\not\subset A$. As homogeneous varieties in projective space are cut out by quadrics (cf [Ram87]), by the Segre embedding we can express $A$ as cut out by varieties of degree $(2,\ldots,2)$ in $\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$.
@@
Suppose now that $\Sigma\not\subset Z$. We now intersect $\Sigma$ and $Z$ to get
@@
(3.3) $\deg Q\geq-\left(\frac{1}{d_{1}}-1\right)(H_{1}\cdot C)-\frac{d_{2}}{d_{1}}(H_{2}\cdot C)-\cdots-\frac{d_{m}}{d_{1}}(H_{m}\cdot C).$
@@
First, we observe that $a_{i}\leq-2$ for all $1\leq i\leq m$. Indeed, if $a_{i}\geq-1$, by Lemma 3.3, the dimension of the space of $\mathbb{P}^{N_{i}}$-lines in $A$ would be $\dim A-a_{i}-3\leq\dim A-2$, but then the lines could not cover $A$, which is not the case for a homogeneous variety. And since $d_{i}\geq D-a_{i}-2$, this implies that $d_{i}\geq 4$ for all $1\leq i\leq m$.
On the other hand, an $n$-dimensional projective manifold $Y$ with an ample divisor $H$ satisfies $K_{Y}\geq-(n+1)H$, with equality when $X$ is isomorphic to $\mathbb{P}^{n}$ (cf [KO73].) Thus, looking at each projection $\pi_{i}:A\to\mathbb{P}^{N_{i}}$, we obtain that $a_{i}\geq-(D+1)$ for each $1\leq i\leq m$.
@@ 4. Lines in hypersurfaces of homogeneous varieties
(cf [Liu22, Proposition 3.1]) Let $d\geq 2$, $m>0$, and let $X\subset G(2,N)$ be a (possibly reducible) variety with class $[X]=m\cdot\sigma_{N-2,N-2-(d+1)}$ in the Chow ring of the Grassmannian of lines $G(2,N)$. Then, the lines parametrized by $X$ pass through a finite set of fixed points in $\mathbb{P}^{N-1}$.
@@
Consider the dual Grassmannian $G(N-2,N)$. The class of the dual variety to $[X]=m\cdot\sigma_{N-2,N-2-(d+1)}$ is $[X^{*}]=m\cdot\sigma_{2^{N-2-(d+1)},1^{d+1}}$.
By Pieri’s formula [3264, Proposition 4.9], $[X^{*}]\cdot\sigma_{2}=0$. Since $\sigma_{2}$ parametrizes $\mathbb{P}^{N-3}$s through a point in $\mathbb{P}^{N-1}$, it follows that the general point of $\mathbb{P}^{N-1}$ is not contained in any of the $\mathbb{P}^{N-3}$s parametrized by $X^{*}$. Hence, the $\mathbb{P}^{N-3}$s parametrized by each component of $X^{*}$ sweep out a subvariety $T$ of dimension $N-2$ of $\mathbb{P}^{N-1}$. We claim that $T$ is a hyperplane, $T\cong\mathbb{P}^{N-2}$. To see this, consider the universal family
$\mathcal{U}=\{(x,p)\ |\ x\in X^{*},\ p\in\mathbb{P}^{N-3}\text{ parametrized by }x\}\subset X^{*}\times T.$
@@
By the degree definition, a general $\mathbb{P}^{N_{1}}$ section of $Z$ is a general degree $d_{1}$ hypersurface in $\mathbb{P}^{N_{1}}$. By [3264, Theorem 6.34], the scheme of lines in a general degree $d_{1}$ hypersurface has codimension $d_{1}+1$ in $G(2,N_{1}+1)$. Thus, $F_{1}^{N_{1}}(Z)$ has codimension $d_{1}+1$ in $G(2,N_{1}+1)\times\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$. On the other hand, by Lemma 3.3, $\dim F_{1}^{N_{1}}(A)=D-a_{1}-3$. Therefore, it suffices to prove the theorem for $D-a_{1}-3=d_{1}+1$, or equivalently, $d_{1}=D-a_{1}-4$.
The intersection of $[F_{1}^{N_{1}}(Z)]$ with $H_{2}^{N_{2}}\cdots H_{m}^{N_{m}}$ is the class of lines in a general $\mathbb{P}^{N_{1}}$ section of $Z$, which is a general degree $d_{1}$ hypersurface in $\mathbb{P}^{N_{1}}$. So, by [3264, Proposition 6.4], we can compute the class $[F_{1}^{N_{1}}(Z)]$:
@@
And since $F_{1}^{N_{1}}(A)$ is a space of $\mathbb{P}^{N_{1}}$-lines, its class $[F_{1}^{N_{1}}(A)]$ is also generated by classes of $G(2,N_{1})$.
By the Whitney formula and the splitting principle [3264, Chapter 5], we can write
@@
Notice that, after expanding terms, all coefficients are non-negative and the term $\sigma_{1,1}\sigma_{1}^{d_{1}-1}$ appears. Recall that these Schubert classes multiply by the rule: $\sigma_{1}\cdot\sigma_{k,l}=\sigma_{k+1,l}+\sigma_{k,l+1}$ and $\sigma_{1,1}\cdot\sigma_{k,l}=\sigma_{k+1,l+1}$. Thus, we can see that in the expansion of $\sigma_{1}^{d_{1}-1}$ all Schubert classes $\sigma_{i,j}$ with $i+j=d_{1}-1$ appear with positive coefficients. It follows that in the expansion of $[F_{1}^{N_{1}}(Z)]$ all classes $\sigma_{i,j}$ with $i+j=d_{1}+1$ and $i,j\geq 1$ appear with positive coefficients, that is, the only degree $d_{1}+1$ class not in the expansion of $[F_{1}^{N_{1}}(Z)]$ is $\sigma_{d_{1}+1,0}$. Since the complementary dimension class of $\sigma_{d_{1}+1,0}$ in $G(2,N+1)$ is $\sigma_{N-1,N-1-(d_{1}+1)}$, it suffices to show that $[F_{1}^{N_{1}}(A)]$ is not a multiple of $\sigma_{N-1,N-1-(d_{1}+1)}$. But, by Lemma 4.1, if $[F_{1}^{N_{1}}(A)]$ is a multiple of $\sigma_{N-1,N-1-(d_{1}+1)}$, then the lines of a general $\mathbb{P}^{N_{1}}$ section of $A$ pass through a finite set of points, what cannot happen since $A$ is homogeneous. Therefore, the product $[F_{1}^{N_{1}}(Z)]\cdot[F_{1}^{N_{1}}(A)]$ is not zero, hence there are $\mathbb{P}^{N_{1}}$-lines contained in $X$.
@@
This leaves open the case $d_{i}=\dim A-a_{i}-3$ for some $1\leq i\leq m$. When $\dim A\leq 4$, there are examples when this bound fails to imply algebraic hyperbolicity:
@@
(cf [Y22, Lemma 4.2]) A very general surface $X\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$ of degree $(d_{1},d_{2})$ with $d_{1}=4$ or $d_{2}=4$ contains an elliptic curve. In particular, it is not algebraically hyperbolic.
@@
A very general surface $X\subset\mathbb{P}^{2}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ of degree $(d_{1},d_{2},d_{3})$ with $d_{1}=4$ contains an elliptic curve. In particular, it is not algebraically hyperbolic.
The bound $d_{i}\geq\dim A-a_{i}-3$ for all $i$ also fails to imply algebraic hyperbolicity in the cases $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\mathbb{P}^{2}\times\mathbb{P}^{1}$ as proved in [CR19]. For $A=\mathbb{P}^{4}$, the case of sextic threefolds remains open.
However, when $\dim A\geq 5$, it does imply algebraic hyperbolicity in the cases $A=\mathbb{P}^{n}$ and $A=\mathbb{P}^{m}\times\mathbb{P}^{n}$ as proved in [Y22]. We conjecture the same holds for homogenous varieties.

### v3

@@ 3. Scroll method for homogeneous varieties
In this section, we work out the scroll method introduced by Coskun and Riedl in [CR23] in a more general setting. This technique builds a surface scroll on the curve $Y_{t}$ whose degree can be used to improve the bound for algebraic hyperbolicity obtained in Proposition 2.9.
Here we will work on homogeneous varieties satisfying the Main Setting 1.3, which we recall here:
Let $A$ be a smooth complex projective variety of dimension $D$ such that:
@@
Let $\mathcal{E}$ be the line bundle $d_{1}H_{1}+\cdots+d_{m}H_{m}$ for $d_{1},\ldots,d_{m}>0$, equivariant under $G$, and assume that $H_{1},\ldots,H_{m}$ is a collection of section-dominating line bundles for $\mathcal{E}$. To simplify the notation, we write $X$ for the zero locus of a very general section of $\mathcal{E}$, and $C$ for the curve in $X$, instead of $X_{t}$ and $Y_{t}$.
@@
are $\mathbb{P}^{N_{1}}$. We define a $\mathbb{P}^{N_{1}}$-line as a line contained in one of these fibers. Equivalently, a $\mathbb{P}^{N_{1}}$-line is a curve of numerical class $H_{1}^{N_{1}-1}H_{2}^{N_{2}}\cdots H_{m}^{N_{m}}$. We define $\mathbb{P}^{N_{i}}$-lines similarly for $i=1,\ldots,m$. We call a surface scroll a $\mathbb{P}^{N_{i}}$-scroll if it contains a $\mathbb{P}^{N_{i}}$-line through every point.
@@
Let $L$ be a $\mathbb{P}^{N_{i}}$-line in $A$. Since $A$ is a homogeneous variety, its tangent bundle $T_{A}$ is globally generated, thus $T_{A}|_{L}$ is globally generated. A classic deformation argument (cf. [Deb01, Theorem 2.6]) then shows that $\dim F_{1}^{N_{i}}(A)=h^{0}(L,N_{L/A})=\dim A-K_{A}\cdot L-3=\dim A-a_{i}-3$. ∎
@@
(cf. [CR23, Lemma 2.13]) A rank one quotient $Q$ of $M_{H_{i}}|_{C}$ induces a $\mathbb{P}^{N_{i}}$-scroll $\Sigma$ over $C$ of $H_{i}$-degree equal to $\deg Q+(H_{i}\cdot C)$.
@@
Without loss of generality, assume $s_{1}\geq\cdots\geq s_{m}$. As before, we can assume $s_{1}+\cdots+s_{m}\leq\operatorname{rk}N_{C/X}=D-2$. We divide the proof into the cases $s_{1}\leq D-3$ and $s_{1}=D-2$.
@@
Case 2. Suppose now instead that $s_{1}=D-2$ is the minimum possible value for $s_{1}$ for which $\beta$ is surjective. So we have $s_{2}=\cdots=s_{m}=0$ and a surjection
@@
Since this is the minimum value for $s_{1}$ such that $\beta$ is surjective, this means that the induced map
@@
has a rank-one image $Q$ (cf. [Cle03, Section 2.2].) By Lemma 3.4, this induces a $\mathbb{P}^{N_{1}}$-scroll $\Sigma\subset\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$ containing $C$ of $H_{1}$-degree $\Sigma\cdot H_{1}^{2}=\deg Q+(H_{1}\cdot C)$. And since $\Sigma$ is a $\mathbb{P}^{N_{1}}$-scroll, we have $H_{1}H_{j}\cdot\Sigma=H_{j}\cdot C$ for $j\neq 1$. Thus, $\Sigma$ has numerical class
@@
Suppose first that $\Sigma\subset Z$. We will prove in Theorem 4.2 that $X$ does not contain $\mathbb{P}^{N_{1}}$-lines. Thus, we must have $\Sigma\not\subset A$. As projective homogeneous varieties in projective space are cut out by quadrics (cf. [Ram87, Theorem 3.11 and the following Remark 3.12]), by the Segre embedding we can express $A$ as cut out by varieties of degree (2,,2)(2$\ldots$,2)( 2 , … , 2 ) in $\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$.
@@
Therefore, in this case we have $2g(C)-2\geq\epsilon\sum_{i=1}^{m}(H_{i}\cdot C)$ for $\epsilon=\frac{1}{2}$, hence $X$ is algebraically hyperbolic.
Suppose now instead that $\Sigma\not\subset Z$. We now intersect $\Sigma$ and $Z$ to get
@@
(3.3) $\deg Q\geq\left(\frac{1}{d_{1}}-1\right)(H_{1}\cdot C)-\frac{d_{2}}{d_{1}}(H_{2}\cdot C)-\cdots-\frac{d_{m}}{d_{1}}(H_{m}\cdot C).$
@@
First, we observe that $a_{i}\leq-2$ for all $1\leq i\leq m$. Indeed, if $a_{i}\geq-1$, by Lemma 3.3, the dimension of the space of $\mathbb{P}^{N_{i}}$-lines in $A$ would be $\dim A-a_{i}-3\leq\dim A-2$, but then the lines could not cover $A$, which is not the case for a homogeneous variety (cf. [Man20, Theorem 1.5 and the following paragraph]). And since $d_{i}\geq D-a_{i}-2$, this implies that $d_{i}\geq 4$ for all $1\leq i\leq m$.
On the other hand, an $n$-dimensional projective manifold $Y$ with an ample divisor $H$ satisfies $K_{Y}\geq-(n+1)H$, with equality when $X$ is isomorphic to $\mathbb{P}^{n}$ (cf. [KO73].) Thus, looking at each projection $\pi_{i}:A\to\mathbb{P}^{N_{i}}$, we obtain that $a_{i}\geq-(D+1)$ for each $1\leq i\leq m$.
@@ 4. Lines in hypersurfaces of homogeneous varieties
(cf. [Liu22, Proposition 3.1]) Let $d\geq 2$, $m>0$, and let $X\subset G(2,N)$ be a (possibly reducible) variety with class $[X]=m\cdot\sigma_{N-2,N-2-(d+1)}$ in the Chow ring of the Grassmannian of lines $G(2,N)$. Then, the lines parametrized by $X$ pass through a finite set of fixed points in $\mathbb{P}^{N-1}$.
@@
Consider the dual Grassmannian $G(N-2,N)$. The class of the dual variety to $[X]=m\cdot\sigma_{N-2,N-2-(d+1)}$ in $G(N-2,N)$ is $[X^{*}]=m\cdot\sigma_{2^{N-2-(d+1)},1^{d+1}}$.
By Pieri’s formula [EH16, Proposition 4.9], $[X^{*}]\cdot\sigma_{2}=0$. Since $\sigma_{2}$ parametrizes $\mathbb{P}^{N-3}$s through a point in $\mathbb{P}^{N-1}$, it follows that the general point of $\mathbb{P}^{N-1}$ is not contained in any of the $\mathbb{P}^{N-3}$s parametrized by $X^{*}$. Hence, the $\mathbb{P}^{N-3}$s parametrized by each component of $X^{*}$ sweep out a subvariety $T$ of dimension $N-2$ of $\mathbb{P}^{N-1}$. We claim that $T$ is a hyperplane, $T\cong\mathbb{P}^{N-2}$. To see this, consider the universal family
$\mathcal{U}=\{(x,p)\ |\ x\in X^{*},\ p\text{ lies in the }\mathbb{P}^{N-3}\text{ parametrized by }x\}\subset X^{*}\times T.$
@@
By the degree definition, a general $\mathbb{P}^{N_{1}}$ section of $Z$ is a general degree $d_{1}$ hypersurface in $\mathbb{P}^{N_{1}}$. By [EH16, Theorem 6.34], the scheme of lines in a general degree $d_{1}$ hypersurface has codimension $d_{1}+1$ in $G(2,N_{1}+1)$. Thus, $F_{1}^{N_{1}}(Z)$ has codimension $d_{1}+1$ in $G(2,N_{1}+1)\times\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$. On the other hand, by Lemma 3.3, $\dim F_{1}^{N_{1}}(A)=D-a_{1}-3$. Therefore, it suffices to prove the theorem for $D-a_{1}-3=d_{1}+1$, or equivalently, $d_{1}=D-a_{1}-4$.
The intersection of $[F_{1}^{N_{1}}(Z)]$ with $H_{2}^{N_{2}}\cdots H_{m}^{N_{m}}$ is the class of lines in a general $\mathbb{P}^{N_{1}}$ section of $Z$, which is a general degree $d_{1}$ hypersurface in $\mathbb{P}^{N_{1}}$. So, by [EH16, Proposition 6.4], we can compute the class $[F_{1}^{N_{1}}(Z)]$:
@@
And since $F_{1}^{N_{1}}(A)$ is a space of $\mathbb{P}^{N_{1}}$-lines, its class $[F_{1}^{N_{1}}(A)]$ is also generated by classes of $G(2,N_{1}+1)$.
By the Whitney formula and the splitting principle [EH16, Chapter 5], we can write
@@
Notice that, after expanding terms, all coefficients are non-negative and the term $\sigma_{1,1}\sigma_{1}^{d_{1}-1}$ appears. Recall that these Schubert classes multiply by the rule: $\sigma_{1}\cdot\sigma_{k,l}=\sigma_{k+1,l}+\sigma_{k,l+1}$ and $\sigma_{1,1}\cdot\sigma_{k,l}=\sigma_{k+1,l+1}$. Thus, we can see that in the expansion of $\sigma_{1}^{d_{1}-1}$ all Schubert classes $\sigma_{i,j}$ with $i+j=d_{1}-1$ appear with positive coefficients. It follows that in the expansion of $[F_{1}^{N_{1}}(Z)]$ all classes $\sigma_{i,j}$ with $i+j=d_{1}+1$ and $i,j\geq 1$ appear with positive coefficients, that is, the only degree $d_{1}+1$ class not in the expansion of $[F_{1}^{N_{1}}(Z)]$ is $\sigma_{d_{1}+1,0}$. Since the complementary dimension class of $\sigma_{d_{1}+1,0}$ in $G(2,N+1)$ is $\sigma_{N-1,N-1-(d_{1}+1)}$, it suffices to show that $[F_{1}^{N_{1}}(A)]$ is not a multiple of $\sigma_{N-1,N-1-(d_{1}+1)}$. But, by Lemma 4.1, if $[F_{1}^{N_{1}}(A)]$ is a multiple of $\sigma_{N-1,N-1-(d_{1}+1)}$, then the lines of a general $\mathbb{P}^{N_{1}}$ section of $A$ pass through a finite set of points, which cannot happen since $A$ is homogeneous. Therefore, the product $[F_{1}^{N_{1}}(Z)]\cdot[F_{1}^{N_{1}}(A)]$ is not zero, hence there are $\mathbb{P}^{N_{1}}$-lines contained in $X$.
@@
  • If $d_{i}\leq\dim A-a_{i}-4$ for some $1\leq i\leq m$, then a general hypersurface $X$ of degree $(d_{1},\ldots d_{m})$ contains $\mathbb{P}^{N_{i}}$-lines. In particular, $X$ is not algebraically hyperbolic.
This leaves open the cases with $d_{i}\geq\dim A-a_{i}-3$ for all $1\leq i\leq m$ and $d_{i}=\dim A-a_{i}-3$ for some $1\leq i\leq m$. When $\dim A\leq 4$, there are examples when this bound fails to imply algebraic hyperbolicity:
@@
(cf. [Yeo25, Lemma 4.2]) A very general hypersurface $X\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$ of degree $(d_{1},d_{2})$ with $d_{1}=4$ or $d_{2}=4$ contains an elliptic curve. In particular, it is not algebraically hyperbolic.
@@
A very general hypersurface $X\subset\mathbb{P}^{2}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ of degree $(d_{1},d_{2},d_{3})$ with $d_{1}=4$ contains an elliptic curve. In particular, it is not algebraically hyperbolic.
The bound $d_{i}\geq\dim A-a_{i}-3$ for all $i$ also fails to imply algebraic hyperbolicity in the cases $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\mathbb{P}^{2}\times\mathbb{P}^{1}$ as proved in [CR23]. For $A=\mathbb{P}^{4}$, the case of sextic threefolds remains open.
However, when $\dim A\geq 5$, it does imply algebraic hyperbolicity in the cases $A=\mathbb{P}^{n}$ and $A=\mathbb{P}^{m}\times\mathbb{P}^{n}$ as proved in [Yeo25]. We conjecture the same holds for homogeneous varieties.